OpenStudy (anonymous):

LaTex doesn't seem to be working with me. :/ I went to welcome a new user, but it wouldn't work. I typed it into the equation editor, waited 'till it loaded in.... and nothing shows up.

OpenStudy (dmezzullo):

@ParthKohli

OpenStudy (swag):

@UnkleRhaukus Should be able to help.

OpenStudy (anonymous):

Can anyone check out my question by any chance?

Parth (parthkohli):

Seems like you're off by a "}" at the end. :-)

OpenStudy (unklerhaukus):

``` \[\Huge{\color{teal}{\mathbb{\text{Welcome To Openstudy!}}}}\] ``` \[\Huge{\color{teal}{\mathbb{\text{Welcome To Openstudy!}}}}\]

Parth (parthkohli):

Oh, that makes two.

OpenStudy (swag):

:)

OpenStudy (anonymous):

Oh wow. lol. \[\Huge{\color{teal}{\mathbb{\text{Thank you☺}}}}\]